OpenStudy (anonymous):

how do you solve 6 3n+6=1296

OpenStudy (callisto):

I assume the question to be:\(6^{3n+6}=1296\)\[6^{3n+6}=1296\]\[6^{3n+6}=6^4\]\[3n+6 = 4\]Solve n :)

OpenStudy (anonymous):

thank you(: